what is the coordinates of incentre
?


 


 

Grade:12

2 Answers

Godfrey Classic Prince
11 years ago

Dear shweta ramki,

The Co-ordinates of the Incentre can be got by using this formula,

Co-ordinates of the Incentre  =   [ (ax1+bx2+cx3)/ (a+b+c) , (ay1+by2+cy3)/ (a+b+c) ]


Side BC=a, Side AB=c, Side AB=c

Distance Formula :

d=\sqrt{(\Delta x)^2+(\Delta y)^2}=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}.\,

 

Hope this helped you immensely..!

All the Very Best & Good Luck to you ..!!

 

Best Regards,

AskIITians Expert,

Godfrey Classic Prince

IIT-M

Please approve my answer if you liked it by clicking on "Yes" given below..!!Smile


 



TANAYRAJ SINGH CHOUHAN
11 years ago

Let there be a triangle ABC with A(x1,y1) , B(x2,y2)  & C(x3,y3)

Let I(h,k) be the incenter of the ΔABC

AND let ''a'' be the side length opposite to A ''b'' opposite to B and ''c'' opposite to C

So the coordinates of I(h,k) = (ax1+bx2+cx3/a+b+c)  ,  (ay1+by2+cy3/a+b+c)

Plz. approve my answer by clicking Yes given below, if u loved it... Plz/....

And best of luck for your preparation!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!